OpenStudy (anonymous):

#1: A material has a density of 8.9 g/cm^3. You have six cubic centimeters of the substance. What is the material’s mass in grams? A. 3.0 g b. 5.9 g C. 27 g D. 53 g **Idk i think it is either C. 27 g or D. 53 g :/

OpenStudy (aaronq):

d=m/V use some math!

OpenStudy (aaronq):

dont guesss haha

OpenStudy (anonymous):

#2: Which of the following describes an element? A. H2O B. CH4 C. H2SO4 D. N2 **I kinda forgot about elements... :/ i think it is D. N2 tho if i remember correctly.. but can u pls explain some more?? :) and is D even correct?? :/

OpenStudy (anonymous):

**sorry my internet got cut off.. but hold on lemme try to solve #1 :)

OpenStudy (anonymous):

8.9 g/cm^3=m/6 right??

OpenStudy (anonymous):

is that right @aaronq ??? and then i multiply each side by 6???? is that correct???

OpenStudy (aaronq):

you're right about #2. it is D, while it's still a molecule (a diatomic one (2 atoms)) it consists of only 1 element, while the others are compounds, meaning at least 2 different atoms, and these are not considered elements.

OpenStudy (anonymous):

so yay!! #2 = D. N2 :) and did i get the steps right for #1 above??

OpenStudy (aaronq):

yes #1 is good (Y)

OpenStudy (anonymous):

8.9 g/cm^3=m/6 multiply both sides by 6 and i get this??? m=53.4 which is about 53. So my answer to #1 is D. 53 g???

OpenStudy (aaronq):

yep! try doing these without a calculator ;)

OpenStudy (anonymous):

okay :) but yayy!! :) thank youuu :D